OpenStudy (anonymous):

@jim_thompson5910 I need so help can you help me please

jimthompson5910 (jim_thompson5910):

with what?

OpenStudy (anonymous):

okay the first thing is how do I multiply this

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

you can use the rule \[\Large \sqrt{x}*\sqrt{y} = \sqrt{x*y}\]

OpenStudy (anonymous):

wat would be the x and y

jimthompson5910 (jim_thompson5910):